OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 45x2 - 196

OpenStudy (anonymous):

Please help :)

OpenStudy (irishboy123):

dividing that function by \(x + 2i\) is not going to be much of a laugh but \(\large f(x) = x^4 - 45x^2 - 196\) \(\large \implies f(p) = p^2 - 45 p -196 = 0 \ \ \ :x = + p^{1/2}\) \(\large \implies p = \frac{45 \pm \sqrt{45^2 - 4(1)(-196)}}{2} = -4 ,49\) so x is what ?!?!

OpenStudy (anonymous):

2i, 14, -14 ??